OpenStudy (jollypikachu300):

WILL FAN AND MEDAL HELP A triangle has a base of 11 inches and a height of 8 inches. What is the area of the triangle? A. 121 square inches B. 64 square inches C. 88 square inches D. 44 square inches

OpenStudy (photon336):

\[Area = \frac{ 1 }{ 2 }base*height\]

OpenStudy (photon336):

plug in the value for base and height \[\frac{ 1 }{ 2 }*11*8 = ? \]

OpenStudy (photon336):

@Jollypikachu300